# Global analysis of charm mixing parameters and determination of the CKM angle $\mathbf{\gamma}$

### v1

@@ III.2 SCS decays to CP eigenstates
where we have neglected the quadratic terms in the charm mixing and CP-violating parameters that come with higher powers in $\tau$. Then, the slope $\tilde{y}_{\text{CP}}$ is extracted from Eq. (24) through a linear fit Aitala et al. (1999); Aaij et al. (2022); Link et al. (2000); Csorna et al. (2002); Zupanc et al. (2009); Lees et al. (2013); Ablikim et al. (2015); Starič et al. (2016); Aaij et al. (2019b); Nayak et al. (2020), and we have:
@@
Experiments measure the slopes $\Delta Y^{f}$ of Eq. (26) for the modes $D^{0}\to K^{+}K^{-}$ and $D^{0}\to\pi^{+}\pi^{-}$aij et al. (2021a), through the following combinations
@@
We use the relations in Eq. (27) to constrain the parameter $\Delta Y^{KK}$, which enters the observables coming from the WS/RS ratios of $D^{0}\to K^{\pm}\pi^{\mp}$ in Eq. (20). Moreover, since in the U-spin symmetric limit the final state-dependent contributions to $\Delta Y^{f}$ are equal and opposite in sign for $K^{+}K^{-}$ and $\pi^{+}\pi^{-}$ Kagan and Silvestrini (2021), averaging the two modes allows $\Delta Y$ to be expressed in the framework of approximate universality as
@@
Notice that we do not include the latest measurement of $\Delta Y^{f}$ from $D^{0}\to\pi^{+}\pi^{-}\pi^{0}$ decays by LHCb Aaij et al. (2024g) in the combination, as it is less precise than the observables in Eq. (27). Moreover, the application of approximate universality in this case is not as good as for $\Delta Y$ since there is no U-spin cancellation of final-state dependent contributions.
We consider in the combination also measurements of the time-integrated CP asymmetry for the mode $K^{+}K^{-}$ Aaij et al. (2023a) which, from Eq. (26), can be written as
@@
where we introduced the so-called average decay time $\left\langle\tau\right\rangle^{f}_{E}$, which depends on the final state $f$ and on the experimental environment $E$.
The first signal of direct CP violation in the charm system Aaij et al. (2019a) was observed in the difference between the asymmetries in Eq. (29) for the modes $K^{+}K^{-}$ and $\pi^{+}\pi^{-}$, which we fit as
$\Delta A^{\text{CP}}_{D}=a_{D}^{KK}-a_{D}^{\pi\pi}+\left\langle\tau\right\rangle^{KK}_{E^{\prime}}\Delta Y^{KK}-\left\langle\tau\right\rangle_{E^{\prime}}^{\pi\pi}\Delta Y^{\pi\pi}.$ (30)
@@ III.3 Multi-body final states
The most precise observables for multi-body final states constraining together the charm mixing and CP-violating parameters come from the Dalitz plot analyses of $D^{0}\to K^{0}_{S}\pi^{+}\pi^{-}$ decays Aaij et al. (2019c, 2021b, 2023b). The reconstruction of the two-dimensional phase space and the decay time is performed, in a model-independent way, by partitioning them into bins, which we indicate as ±$\pm i$ and $j$, respectively, with $i$ and $-i$ connected by a CP transformation. For each of the bins, ratios analogous to WS/RS in Eq. (15) are measured:
@@
Besides the Dalitz plot analysis, the phase space dependence of the $D$ mesons decay amplitudes to a multi-body final state $f$, such as $K^{0}_{S}K^{-}\pi^{+}$, can be parametrized through the following two quantities:
@@
where direct CP violation has been neglected. The integrals in Eq. (33) are performed over the $D$ phase space coordinates $\Phi_{D}$. The parameter in the upper row is a straightforward generalization for multi-body final states of the ratios introduced in Eqs. (14) and (23) for two-body decays. The interference integral in the bottom row is described simply through its magnitude $\kappa_{D}^{f}$, also known as the coherence factor, and its integrated strong phase $\delta_{D}^{f}$. For quasi-CP eigenstates, such as $\pi^{+}\pi^{-}\pi^{0}$, the imaginary part of the interference integral in Eq. (33) is vanishing. Then, instead of the coherence factor and the integrated strong phase, it is customary to use one single real parameter: the CP-even fraction $F_{+}^{f}$, defined as
@@
which has already been introduced in Eq. (22). From its definition in Eq. (34), the CP-even fraction is non-negative and it is unity for two-body states.
The decay parameters for $K^{0}_{S}K^{\pm}\pi^{\mp}$ final states are measured from the ratio between the branching fractions of the DCS and CF decays Insler et al. (2012); Aaij et al. (2016b), given by
$\frac{\mathcal{B}(D^{0}\to K^{0}_{S}K^{+}\pi^{-})}{\mathcal{B}(D^{0}\to K^{0}_{S}K^{-}\pi^{+})}=\frac{(r_{D}^{K^{0}_{S}K\pi})^{2}-\kappa_{D}^{K^{0}_{S}K\pi}r_{D}^{K^{0}_{S}K\pi}(y\cos\delta_{D}^{K^{0}_{S}K\pi}-x\sin\delta_{D}^{K^{0}_{S}K\pi})}{1-\kappa_{D}^{K^{0}_{S}K\pi}r_{D}^{K^{0}_{S}K\pi}(y\cos\delta_{D}^{K^{0}_{S}K\pi}+x\sin\delta_{D}^{K^{0}_{S}K\pi})},$ (35)
@@ IV Beauty observables
$A_{B}^{\text{CP}}(XX\pi^{0},h)$, $A^{\text{fav,sup}}(K\pi\pi^{0},h)$, $B^{\pm}\to Dh^{\pm}$ $D\to K\pi\pi^{0}$, Aaij et al. (2015a)
$R^{\text{CP}}(XX\pi^{0},K\pi\pi^{0},K,\pi)$ $R^{\text{ADS}}(K\pi\pi^{0},h)$ $D\to XX\pi^{0}$
$A_{B}^{\text{CP}}(XX,h\pi\pi)$, $A^{\text{fav}}(K\pi,h\pi\pi)$, $B^{\pm}\to Dh^{\pm}\pi\pi$ $D\to K\pi$, Aaij et al. (2015b)
$A_{B}^{\text{CP}}(4\pi,h)$, $A^{\text{fav,sup}}(K3\pi,h)$, $B^{\pm}\to Dh^{\pm}$ $D\to K3\pi$, Aaij et al. (2016d)
$R^{\text{CP}}(4\pi,K3\pi,K,\pi)$ $R^{\text{ADS}}(K3\pi,h)$ $D\to 4\pi$
$A_{B}^{\text{CP}}(XX,K^{*\pm})$, $R(XX,K\pi,K^{*\pm})$ $A^{\text{fav,sup}}(K\pi,K^{*\pm})$, $R^{\text{ADS}}(K\pi,K^{*\pm})$ $B^{\pm}\to DK^{*\pm}$ $D\to K\pi$, $D\to XX$, Aaij et al. (2024c)
$A_{B}^{\text{CP}}(XX,K^{*0})$, $R(XX,K\pi,K^{*0})$ $A^{\text{fav}}(K\pi,K^{*0})$, $R_{\pm}(K\pi,K^{*0})$ $B^{0}\to DK^{*0}$ $D\to K\pi$, $D\to XX$, Aaij et al. (2019d),
$A_{B}^{\text{CP}}(4\pi,K^{*0})$, $R(4\pi,K3\pi,K^{*0})$ $A^{\text{fav}}(K3\pi,K^{*0})$, $R_{\pm}(K3\pi,K^{*0})$ $B^{0}_{s}\to D\overline{K^{*0}}$ $D\to K3\pi$, $D\to 4\pi$ Aaij et al. (2024b)
$A_{B}^{\text{CP}}(XX,h)$, $R^{\text{CP}}(XX,K\pi,K,\pi)$ $A^{\text{fav}}(K\pi,K)$, $R_{\pm}(K\pi,h)$ $B^{\pm}\to Dh^{\pm}$ $D\to K\pi$, $D\to XX$ Aaij et al. (2021c)
$A_{B}^{\text{CP}}(XX,h(\gamma,\pi^{0}))$, $A^{\text{fav}}(K\pi,K(\gamma,\pi^{0}))$, $B^{\pm}\to[D\pi^{0}]_{D^{*}}h^{\pm}$ $D\to K\pi$, Aaij et al. (2021c)
$A^{\text{fav,sup}}(K^{0}_{S}K\pi,h)$,
—— $R^{\text{fav,sup}}(K^{0}_{S}K\pi,K,\pi)$, $B^{\pm}\to Dh^{\pm}$ $D\to K^{0}_{S}K\pi$ Aaij et al. (2020),
$R^{\text{ADS}}(K^{0}_{S}K\pi,\pi)$ Adachi et al. (2023)
$A_{B}^{\text{CP}}(XX\pi\pi,h)$, $R^{\text{CP}}(XX\pi\pi,K3\pi,K,\pi)$ —— $B^{\pm}\to Dh^{\pm}$ $D\to XX\pi\pi$, $D\to K3\pi$ Aaij et al. (2023c)
$R^{\text{ADS}}(K\pi,h)$, $A^{\text{sup}}(K\pi,h)$ $B^{\pm}\to Dh^{\pm}$ $D\to K\pi$ Horii et al. (2011)
$R^{\text{ADS}}(K\pi\pi^{0},h)$, $A^{\text{sup}}(K\pi\pi^{0},h)$ $B^{\pm}\to Dh^{\pm}$ $D\to K\pi\pi^{0}$ Nayak et al. (2013)
$A_{B}^{\text{CP}}(K^{+}K^{-},K)$, $R^{\text{CP}}(K^{+}K^{-},K\pi,K,\pi)$ $B^{\pm}\to Dh^{\pm}$ $D\to K^{+}K^{-}$, $D\to K\pi$, Adachi et al. (2024a)
$A_{B}^{\text{CP}}(f^{\pm},K)$, $R^{\text{CP}}(f^{\pm},K\pi,K,\pi)$ $B^{\pm}\to Dh^{\pm}$ $D\to f^{\pm}$, Abe et al. (2006)
$A_{B}^{\text{CP}}(f^{\pm},K(\pi^{0}))$, —— $B^{\pm}\to[D\pi^{0}]_{D^{*}}h^{\pm}$ $D\to K\pi$
$R^{\text{CP}}(f^{\pm},K\pi,K(\pi^{0}),\pi(\pi^{0}))$
Table 2: GLW and ADS observables coming from $B$ mesons cascade decays. The $XX$ notation indicate both $K^{+}K^{-}$ and $\pi^{+}\pi^{-}$ states, while we use $K\pi$ for $K^{\mp}\pi^{\pm}$. $h$ stands for both $K$ and $\pi$ mesons. We refer to $\pi^{+}\pi^{-}\pi^{+}\pi^{-}$ and $K^{\mp}\pi^{\pm}\pi^{+}\pi^{-}$ final states as $4\pi$ and $K3\pi$, respectively. We use $f^{\pm}$ as a shorthand notation for CP-even and CP-odd eigenstates, respectively. The observables are described in details in Apps. A and B.
The precision on the charm mixing and decay parameters can be improved by including in the combination time-integrated measurements of the so-called $B$ meson cascade decays to $D^{0}-\overline{D^{0}}$ mixed states Rama (2014); Aaij et al. (2021f); LHC (2024). We show how the CKM angle $\gamma$ appears in the equations describing the most used and precise observables available to date.
We discuss the extraction of $\gamma$ from the CP-violating phase of the interference between neutral $B$ decays to charmed mesons with and without mixing that is used to test the consistency of the combination.
@@ IV.1 Cascade decays
Beauty cascade decays Abe et al. (2006); Horii et al. (2011); Nayak et al. (2013); Aaij et al. (2015a, b, 2016d); LHC (2016); Aaij et al. (2016e, 2019d, 2021c, 2020); Adachi et al. (2023, 2024a); Aaij et al. (2024b, c) are processes where a $B$ meson goes to an hadron state $h$ and a $D^{0}$ or $\overline{D^{0}}$ meson, which subsequently oscillates and decays to one of the final states $f$ introduced in Sec. III. In the following, we indicate with $B$ a negatively charged $B^{-}$ or a neutral meson $\overline{B^{0}_{(s)}}$.
@@
$(r_{B}^{Dh})^{2}=\frac{\int\text{d}\Phi_{B}\left|\mathcal{A}_{B}^{\overline{D}h}\right|^{2}}{\int\text{d}\Phi_{B}\left|\mathcal{A}_{B}^{Dh}\right|^{2}},\qquad\qquad\kappa_{B}^{Dh}e^{i(\delta_{B}^{Dh}-\gamma)}=\frac{\int\text{d}\Phi_{B}\mathcal{A}_{B}^{\overline{D}h}(\mathcal{A}_{B}^{Dh})^{*}}{\sqrt{\int\text{d}\Phi_{B}\left|\mathcal{A}_{B}^{Dh}\right|^{2}}\sqrt{\int\text{d}\Phi_{B}\left|\mathcal{A}_{B}^{\overline{D}h}\right|^{2}}},$ (36)
@@
Then, the time-integrated rate of the cascade decay $B\to[f]_{D}h$ is found to be
$\Gamma(B\to[f]_{D}h)\propto 1+(r^{f}_{D}r_{B}^{Dh})^{2}+2\kappa_{D}^{f}\kappa_{B}^{Dh}r_{D}^{f}r_{B}^{Dh}\cos(\delta_{B}^{Dh}-\delta_{D}^{f}-\gamma)+\Gamma_{\mathrm{mix}}(x,y,\gamma,f,h),$ (37)
@@
GGSZ Observables $\mathbf{B}\to\mathbf{D^{0}}$ decays $\mathbf{D^{0}}$ decays Ref.
$x_{\pm}^{DK^{*0}}$, $y_{\pm}^{DK^{*0}}$ $B^{0}\to DK^{*0}$ $D\to K^{0}_{S}XX$ Aaij et al. (2016f, 2024e)
$x^{DK}_{\pm}$, $x^{D\pi}_{\xi}$ $B^{\pm}\to DK^{\pm}$ $D\to K^{0}_{S}XX$ Resmi et al. (2019); Aaij et al. (2021d); Abudinén et al. (2022)

### v3

@@ III.2 SCS decays to CP eigenstates
where we have neglected the quadratic terms in the charm mixing and CP-violating parameters that come with higher powers in $\tau$. Then, the slope $\tilde{y}_{\text{CP}}$ is extracted from Eq. (24) through a linear fit E791:1999bzz ; LHCb:2022gnc ; FOCUS:2000kxx ; CLEO:2001lgl ; Belle:2009xzl ; BaBar:2012bho ; BESIII:2015ado ; Belle:2015etc ; LHCb:2018zpj ; Belle:2019xha , and we have:
@@
Experiments measure the slopes $\Delta Y^{f}$ of Eq. (26) for the modes $D^{0}\to K^{+}K^{-}$ and $D^{0}\to\pi^{+}\pi^{-}$:2014wyb ; LHCb:2015xyd ; LHCb:2017ejh ; LHCb:2019dom ; LHCb:2021vmn ; LHCb:2021vmn , and we can define the following combinations
@@
We use the relations in Eq. (27) to constrain the parameter $\Delta Y^{KK}$, which enters the observables coming from the WS/RS ratios of $D^{0}\to K^{\pm}\pi^{\mp}$ in Eq. (20). Moreover, since in the U-spin symmetric limit the final state-dependent contributions to $\Delta Y^{f}$ are equal and opposite in sign for $K^{+}K^{-}$ and $\pi^{+}\pi^{-}$ Kagan:2020vri , averaging the two modes allows $\Delta Y$ to be expressed in the framework of approximate universality as
@@
Notice that we do not include the latest measurement of $\Delta Y^{f}$ from $D^{0}\to\pi^{+}\pi^{-}\pi^{0}$ decays by LHCb LHCb:2024jpt in the combination, as it is less precise than the observables in Eq. (27). Moreover, the application of approximate universality in this case is not as good as for $\Delta Y$ since there is no U-spin cancellation of final-state dependent contributions.
We consider in the combination also measurements of the time-integrated CP asymmetry for the mode $K^{+}K^{-}$ LHCb:2014kcb ; LHCb:2016nxk ; LHCb:2022lry ; BaBar:2007tfw ; Belle:2008ddg ; DiCanto:2012ufu ; CDF:2011ejf which, from Eq. (26), can be written as
@@
where we introduced the so-called average decay time $\left\langle\tau\right\rangle^{f}_{E}$, which depends on the final state $f$ and on the experimental environment $E$. We assume $\left\langle\tau\right\rangle^{\pi\pi}_{B\text{facts}}=$ $\left\langle\tau\right\rangle^{KK}_{B\text{facts}}=1$ for measurements performed at the $B$ factories.
The first signal of direct CP violation in the charm system LHCb:2019hro was observed in the difference between the asymmetries in Eq. (29) for the modes $K^{+}K^{-}$ and $\pi^{+}\pi^{-}$, which we fit as
$\Delta A^{\text{CP}}_{D}=a_{D}^{KK}-a_{D}^{\pi\pi}+\left\langle\tau\right\rangle^{KK}_{E}\Delta Y^{KK}-\left\langle\tau\right\rangle_{E}^{\pi\pi}\Delta Y^{\pi\pi}.$ (30)
@@ III.3 Multi-body final states
The most precise observables for multi-body final states constraining together the charm mixing and CP-violating parameters come from the Dalitz plot analyses of $D^{0}\to K^{0}_{S}\pi^{+}\pi^{-}$ decays LHCb:2019mxy ; LHCb:2022cak . The reconstruction of the two-dimensional phase space and the decay time is performed, in a model-independent way, by partitioning them into bins, which we indicate as ±$\pm i$ and $j$, respectively, with $i$ and $-i$ connected by a CP transformation. For each of the bins, ratios analogous to WS/RS in Eq. (15) are measured:
@@
Notice that for $D\to K^{0}_{S}\pi\pi$ measurements performed at $B$ factories we take into account $K-\overline{K}$ mixing following Ref. Kagan:2020vri (see Eq. (144)). Besides the Dalitz plot analysis, the phase space dependence of the $D$ mesons decay amplitudes to a multi-body final state $f$, such as $K^{0}_{S}K^{-}\pi^{+}$, can be parametrized through the following two quantities:
@@
where direct CP violation has been neglected. The integrals in Eq. (33) are performed over the $D$ phase space coordinates $\Phi_{D}$. The parameter in the upper row is a straightforward generalization for multi-body final states of the ratios introduced in Eqs. (14) and (23) for two-body decays. The interference integral in the bottom row is described simply through its magnitude $\kappa_{D}^{f}$, also known as the coherence factor, and its integrated strong phase $\delta_{D}^{f}$. For quasi-CP eigenstates, such as $\pi^{+}\pi^{-}\pi^{0}$, the imaginary part of the interference integral over the full phase space in Eq. (33) is vanishing. Then, instead of the coherence factor and the integrated strong phase, it is customary to use one single real parameter: the CP-even fraction $F_{+}^{f}$, defined as
@@
which has already been introduced in Eq. (22). From its definition in Eq. (34), the CP-even fraction is non-negative and it is unity for two-body CP-even states.
The decay parameters for $K^{0}_{S}K^{\pm}\pi^{\mp},\ K^{\pm}\pi^{\mp},$ and $K^{\pm}\pi^{\mp}\pi^{0}$ final states are measured from the ratio between the branching fractions of the DCS and CF decays CLEO:2012obf ; LHCb:2015lnk ; BESIII:2025vbt , given by
$\frac{\mathcal{B}(D^{0}\to\overline{f})}{\mathcal{B}(D^{0}\to f)}=\frac{(r_{D}^{f})^{2}-\kappa_{D}^{f}r_{D}^{f}(y\cos\delta_{D}^{f}-x\sin\delta_{D}^{f})}{1-\kappa_{D}^{f}r_{D}^{f}(y\cos\delta_{D}^{f}+x\sin\delta_{D}^{f})},$ (35)
@@ IV Beauty observables
Several charm mixing and decay parameters also enter time-integrated measurements of the so-called $B$ meson cascade decays to $D^{0}-\overline{D^{0}}$ mixed states Rama:2013voa ; LHCb:2021dcr ; LHCb-CONF-2024-004 . We show how the CKM angle $\gamma$ appears in the equations describing the most used and precise observables available to date. We also discuss the extraction of $\gamma$ from the CP-violating phase of the interference between neutral $B$ decays to charmed mesons with and without mixing.
@@ IV.1 Cascade decays
Beauty cascade decays Belle:2006cuz ; Belle:2011ac ; Belle:2013dtr ; LHCb:2015dlc ; LHCb:2017egy ; LHCb:2016bsl ; LHCb:2020hdx ; LHCb:2020vut ; LHCb:2021mmv ; Belle:2023lha ; Belle:2023yoe ; LHCb:2024oco ; LHCb:2024ett ; LHCb:2022nng ; Belle:2010xyn ; Belle:2019uav ; LHCb:2020yot ; Belle:2021efh ; LHCb:2023yjo ; LHCb:2023lib ; LHCb:2023kpr ; LHCb:2023ayf ; BaBar:2007dro ; BaBar:2008qcq ; BaBar:2009dzx ; BaBar:2010hvw ; BaBar:2010otv ; BaBar:2011rud ; CDF:2009wnr ; CDF:2011xrp ; Belle:2006lys ; BaBar:2010uep ; Belle:2015roy are processes where a $B$ meson goes to an hadron state $h$ and a $D^{0}$ or $\overline{D^{0}}$ meson, which subsequently oscillates and decays to one of the final states $f$ introduced in Sec. III. In the following, we indicate with $B$ a positively charged $B^{+}$ or a neutral meson $B^{0}_{(s)}$.
@@
$(r_{B}^{Dh})^{2}=\frac{\int\text{d}\Phi_{B}\left|\mathcal{A}_{B}^{Dh}\right|^{2}}{\int\text{d}\Phi_{B}\left|\mathcal{A}_{B}^{\overline{D}h}\right|^{2}},\qquad\qquad\kappa_{B}^{Dh}e^{-i(\delta_{B}^{Dh}+\gamma)}=\frac{\int\text{d}\Phi_{B}\mathcal{A}_{B}^{\overline{D}h}(\mathcal{A}_{B}^{Dh})^{*}}{\sqrt{\int\text{d}\Phi_{B}\left|\mathcal{A}_{B}^{Dh}\right|^{2}}\sqrt{\int\text{d}\Phi_{B}\left|\mathcal{A}_{B}^{\overline{D}h}\right|^{2}}},$ (36)
@@
Then, the time-integrated rate of the cascade decay $\overline{B}\to[f]_{D}h$ is found to be
$\Gamma(\overline{B}\to[f]_{D}h)\propto 1+(r^{f}_{D}r_{B}^{Dh})^{2}+2\kappa_{D}^{f}\kappa_{B}^{Dh}r_{D}^{f}r_{B}^{Dh}\cos(\delta_{B}^{Dh}-\delta_{D}^{f}-\gamma)+\Gamma_{\mathrm{mix}}(x,y,\gamma,f,h),$ (37)
@@
The $\alpha$ coefficient in Eq. (38) describes the impact on $\Gamma_{\mathrm{mix}}(x,y,\gamma,f,h)$ of the non-trivial dependence of the signal selection efficiency on time and it is one when the latter is constant Rama:2013voa .
Since at this order in the charm mixing and CP-violating parameters the decay rate of the CP conjugated process is obtained simply by replacing $\gamma$ with $-\gamma$ in Eq. (37), the difference between $\Gamma(\overline{B}\to[f]_{D}h)$ and $\Gamma(B\to[\overline{f}]_{D}\overline{h})$ gives access to $\sin\gamma$. This is commonly referred to as the time-integrated CP asymmetry, which is an example of the so-called Gronau-London-Wyler (GLW) Gronau:1991dp ; Gronau:1990ra and Atwood-Dunietz-Soni (ADS) Atwood:1996ci ; Atwood:2000ck observables that we use in our combination. A complete list of the GLW and ADS modes considered in this work is reported in Tab. 2, while their parametrizations are derived in Apps. A and B. We have taken the coherence factors for the modes $B\to DK^{*}$ from LHCb:2017egy and LHCb:2016bsl for charged and neutral $B$ mesons, respectively. For cascade decays with multi-body final states of the charm Giri:2003ty ; Belle:2004bbr ; BPGGSZ_Belle , as $K^{0}_{S}K^{+}K^{-}$, $K^{0}_{S}\pi^{+}\pi^{-}$, $\pi^{+}\pi^{-}K^{+}K^{-}$, and $\pi^{+}\pi^{-}K^{+}K^{-}$, analyses of the variation of the decay rates across the phase space are available Belle:2019uav ; LHCb:2020yot ; Belle:2021efh ; LHCb:2023yjo ; LHCb:2023lib ; LHCb:2023kpr ; LHCb:2023ayf ; Belle:2010xyn ; LHCb:2024ett . The Dalitz plots of these events are fitted using a model for the $D$ meson decay amplitudes or by binning the phase space and solving a system of linear equations in a model-independent way. The first method is more precise statistically with respect to the latter but introduces an additional systematic uncertainty due to the model.
$A_{B}^{\text{CP}}(XX\pi^{0},h)$, $A^{\text{fav,sup}}(K\pi\pi^{0},h)$, $B^{\pm}\to Dh^{\pm}$ $D\to K\pi\pi^{0}$, LHCb:2021mmv
$R^{\text{CP}}(XX\pi^{0},K\pi\pi^{0},K,\pi)$ $R_{\pm}(K\pi\pi^{0},h)$ $D\to XX\pi^{0}$
$A_{B}^{\text{CP}}(XX,h\pi\pi)$, $A^{\text{fav}}(K\pi,h\pi\pi)$, $B^{\pm}\to Dh^{\pm}\pi\pi$ $D\to K\pi$, LHCb:2015dlc
$A_{B}^{\text{CP}}(XX,K^{*\pm})$, $R(XX,K\pi,K^{*\pm})$ $A^{\text{fav,sup}}(K\pi,K^{*\pm})$, $R^{\text{ADS}}(K\pi,K^{*\pm})$ $B^{\pm}\to DK^{*\pm}$ $D\to K\pi$, $D\to XX$, LHCb:2024ett
$A_{B}^{\text{CP}}(XX,K^{*0})$, $R(XX,K\pi,K^{*0})$ $A^{\text{fav}}(K\pi,K^{*0})$, $R_{\pm}(K\pi,K^{*0})$ $B^{0}\to DK^{*0}$ $D\to K\pi$, $D\to XX$, LHCb:2024oco
$A_{B}^{\text{CP}}(4\pi,K^{*0})$, $R(4\pi,K3\pi,K^{*0})$ $A^{\text{fav}}(K3\pi,K^{*0})$, $R_{\pm}(K3\pi,K^{*0})$ $B^{0}_{s}\to D\overline{K^{*0}}$ $D\to K3\pi$, $D\to 4\pi$
$A_{B}^{\text{CP}}(XX,h)$, $R^{\text{CP}}(XX,K\pi,K,\pi)$ $A^{\text{fav}}(K\pi,K)$, $R_{\pm}(K\pi,h)$ $B^{\pm}\to Dh^{\pm}$ $D\to K\pi$, $D\to XX$ LHCb:2020hdx
$A_{B}^{\text{CP}}(XX,h(\gamma,\pi^{0}))$, $A^{\text{fav}}(K\pi,K(\gamma,\pi^{0}))$, $B^{\pm}\to[D\pi^{0}]_{D^{*}}h^{\pm}$ $D\to K\pi$, LHCb:2020hdx
$A^{\text{fav,sup}}(K^{0}_{S}K\pi,h)$,
—— $R^{\text{fav,sup}}(K^{0}_{S}K\pi,K,\pi)$, $B^{\pm}\to Dh^{\pm}$ $D\to K^{0}_{S}K\pi$ LHCb:2020vut ,
$R^{\text{ADS}}(K^{0}_{S}K\pi,\pi)$ Belle:2023lha
$A_{B}^{\text{CP}}(XX\pi\pi,h)$, $R^{\text{CP}}(XX\pi\pi,K3\pi,K,\pi)$ —— $B^{\pm}\to Dh^{\pm}$ $D\to XX\pi\pi$ LHCb:2023yjo
—— $R^{\text{ADS}}(K\pi,h)$, $A^{\text{sup}}(K\pi,h)$ $B^{\pm}\to Dh^{\pm}$ $D\to K\pi$ Belle:2011ac
—— $R^{\text{ADS}}(K\pi\pi^{0},h)$, $A^{\text{sup}}(K\pi\pi^{0},h)$ $B^{\pm}\to Dh^{\pm}$ $D\to K\pi\pi^{0}$ Belle:2013dtr
$A_{B}^{\text{CP}}(K^{+}K^{-},K)$, $R^{\text{CP}}(K^{+}K^{-},K\pi,K,\pi)$ $B^{\pm}\to Dh^{\pm}$ $D\to K^{+}K^{-}$, $D\to K\pi$, Belle:2023yoe
$A_{B}^{\text{CP}}(f^{\pm},K(\pi^{0}))$, $B^{\pm}\to[D\pi^{0}]_{D^{*}}h^{\pm}$ $D\to f^{\pm}$, Belle:2006cuz
$R^{\text{CP}}(f^{\pm},K\pi,K(\pi^{0}),\pi(\pi^{0}))$ —— $D\to K\pi$
$A_{B}^{\text{CP}}(f^{\pm},K)$, $R^{\text{CP}}(f^{\pm},K\pi,K,\pi)$ —— $B^{\pm}\to DK^{\pm}$ $D\to f^{\pm}$, $D\to K\pi$ BaBar:2010hvw
$A_{B}^{\text{CP}}(f^{\pm},K)$, $R^{\text{CP}}(f^{\pm},K\pi,K,\pi)$ —— $B^{\pm}\to D^{*}K^{\pm}$ $D\to f^{\pm}$, $D\to K\pi$ BaBar:2008qcq
$A_{B}^{\text{CP}}(f^{\pm},K^{*\pm})$, $R(f^{\pm},K\pi,K^{*\pm})$ $A^{\mathrm{sup}}(K\pi,K^{*\pm})$, $R^{\mathrm{ADS}}(K\pi,K^{*\pm})$ $B^{\pm}\to DK^{*\pm}$ $D\to f^{\pm}$, $D\to K\pi$ BaBar:2009dzx
$A_{B}^{\text{CP}}(\pi^{+}\pi^{-}\pi^{0},K)$ —— $B^{\pm}\to DK^{\pm}$ $D\to\pi^{+}\pi^{-}\pi^{0}$ BaBar:2007dro
$[R_{-}(K\pi,h(\gamma,\pi^{0}))+R_{+}(K\pi,h(\gamma,\pi^{0}))]/2$ $B^{\pm}\to[D\pi^{0}]_{D^{*}}h^{\pm}$ $D\to K\pi$ BaBar:2010otv
—— $\frac{R_{-}(K\pi,h(\gamma,\pi^{0}))-R_{+}(K\pi,h(\gamma,\pi^{0}))}{R_{-}(K\pi,h(\gamma,\pi^{0}))+R_{+}(K\pi,h(\gamma,\pi^{0}))}$ $B^{\pm}\to[D\gamma]_{D^{*}}h^{\pm}$
$[R_{-}(K\pi,h)+R_{+}(K\pi,h)]/2$ $B^{\pm}\to Dh^{\pm}$ $D\to K\pi$ BaBar:2010otv
—— $\frac{R_{-}(K\pi,h)-R_{+}(K\pi,h)}{R_{-}(K\pi,h)+R_{+}(K\pi,h)}$
—— $R_{\pm}(K\pi\pi^{0},K)$ $B^{\pm}\to DK^{\pm}$ $D\to K\pi\pi^{0}$ BaBar:2011rud
$A_{B}^{\text{CP}}(XX,K)$, $R^{\text{CP}}(XX,K\pi,K,\pi)$ —— $B^{\pm}\to DK^{\pm}$ $D\to XX$, $D\to K\pi$ CDF:2009wnr
—— $A^{\mathrm{sup}}(K\pi,h)$, $R^{\mathrm{ADS}}(K\pi,h)$ $B^{\pm}\to Dh^{\pm}$ $D\to K\pi$ CDF:2011xrp
Table 2: GLW and ADS observables coming from $B$ mesons cascade decays. The $XX$ notation denotes both $K^{+}K^{-}$ and $\pi^{+}\pi^{-}$ states, while we use $K\pi$ for $K^{\mp}\pi^{\pm}$. $h$ stands for both $K$ and $\pi$ mesons. We refer to $\pi^{+}\pi^{-}\pi^{+}\pi^{-}$ and $K^{\mp}\pi^{\pm}\pi^{+}\pi^{-}$ final states as $4\pi$ and $K3\pi$, respectively. We use $f^{\pm}$ as a shorthand notation for CP-even and CP-odd eigenstates, respectively. The observables are described in details in Apps. A and B.
In both cases, the following cartesian observables are obtained:
Measuring the quantities in Eq. (39) for beauty decays with a pion in the final state is particularly hard due to the small value of $r_{B}^{D\pi}\sim\mathcal{O}(10^{-3})$.
BPGGSZ cartesian observables $\mathbf{B}\to\mathbf{D^{0}}$ decays $\mathbf{D^{0}}$ decays Ref.
$x_{\pm}^{DK^{*0}}$, $y_{\pm}^{DK^{*0}}$ $B^{0}\to DK^{*0}$ $D\to K^{0}_{S}XX$ LHCb:2023ayf
$x^{DK}_{\pm}$, $x^{D\pi}_{\xi}$ $B^{\pm}\to DK^{\pm}$ $D\to K^{0}_{S}\pi^{+}\pi^{-}\pi^{0}$ Belle:2019uav